# A Neural Material Point Method for Particle-based Simulations

### v1

@@ 4 Experiments
We conduct a set of diverse experiments to demonstrate the accuracy, speed, and generalization capabilities of NeuralMPM. Specifically, we examine its robustness to hyperparameter and architectural choices through an ablation study (4.1). We compare NeuralMPM to GNS [18] and DMCF [20] in terms of accuracy, training time, convergence, and inference speed (4.2). We also evaluate the generalization capabilities of NeuralMPM (4.3) and illustrate how its differentiability can be leveraged to solve an inverse design problem (4.4). Through these experiments, we demonstrate that NeuralMPM is a flexible, accurate, and fast method for emulating complex particle-based simulations.
@@
We consider 6 datasets with variable sequence lengths, numbers of particles, and materials. The first three datasets, WaterRamps, SandRamps, and Goop, contain a single material, water, sand, and goop, respectively, with different material properties. The first two datasets contain random ramp obstacles to challenge the model’s generalization capacity. The fourth dataset, MultiMaterial, mixes the three materials together in the same simulations. These four datasets are taken from [18] and were simulated using the Taichi-MPM simulator [52]. They each contain 1000 trajectories for training and 30 (Goop) or 100 (WaterRamps, SandRamps, MultiMaterial) for validation and testing. The fifth dataset, WBC-SPH [20], was generated using a high-fidelity SPH solver [53] and contains water, random obstacles, and variable gravity. It contains 30 trajectories for training and 9 for validation and testing. The last dataset, Dam Break 2D, was also generated using SPH and contains 50 trajectories for learning, and 25 for validation and testing. Rollout snapshots of NeuralMPM compared to the ground truth for each dataset are shown in Figure 2.
@@
NeuralMPM is trained on a set of full trajectories, with varying initial conditions and number of particles. The training batches are sampled randomly in time and across sequences. We use the Adam optimizer [54] with the following learning rate schedule: a linear warm-up over 100 steps from $10^{-5}$ to $10^{-3}$, $900$ steps at $10^{-3}$, then a cosine annealing [55] for $100,000$ iterations. We use a batch size of $128$, $K=4$ autoregressive steps per iteration, bundle $m=8$ timesteps per model call (resulting in $24$ predicted states), and a grid size of $64\times 64$. For most of our experiments, we use a U-Net [50] with three downsampling blocks with a factor of $2$, $64$ hidden channels, a kernel size of $3$, and MLPs with three hidden layers of size 66664 for pixel-wise encoding and decoding into a latent space. For a fair comparison, we ran training and inference for NeuralMPM, DMCF, and GNS on the exact same hardware. GNS and DMCF were trained until convergence (a maximum of 120 and 240 hours, respectively), while NeuralMPM required 20 hours or less to converge. Further details on training can be found in Appendix A.
WaterRamps SandRamps Goop MultiMaterial WBC-SPH Dam Break 2D
Initial
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Predictions
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ground Truth
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Figure 2: Example snapshots. We train and evaluate NeuralMPM on WaterRamps, SandRamps and Goop, each consisting of a single material, on MultiMaterial that mixes water, sand and goop, on WBC-SPH with more complex obstacles and variable gravity, and on Dam Break 2D, a rectangular-shaped SPH dataset. NeuralMPM is able to learn various kinds of materials, their interactions, and their interactions with solid obstacles. Despite being inspired by MPM, it is not limited to data showing MPM-like behaviour.
@@
Figure 3: Ablation results. Mean squared error (MSE) of full rollouts on unseen test data for Goop. The default parameters are in blue. The dashed orange line shows the best rollout MSE of GNS as reported in [18] on the same data ($1.89\times 10^{-3}$), while the dotted line ($2.4\times 10^{-3}$) indicates the MSE we obtained for GNS after $240$ hours ($20$M training steps). NeuralMPM is robust to hyperparameter changes, with the biggest effects coming from the number of timesteps bundled together ($m$) and grid noise. For a rollout of length $T$, the model is called $T/m$ times, meaning lower values of $m$ require maintaining stability for longer. Autoregressive training coupled with time bundling is sufficient to stabilize the model, eliminating the need for noise injection. Although GNS reportedly outperforms NeuralMPM by a small margin, these results could not be reproduced in our experiments.
@@ 4.1 Ablation study
Figure 3 summarizes the ablation results. A larger number $K$ of autoregressive steps yields more accurate rollouts without the need to add noise. Indeed, injecting noise does not improve accuracy and is even detrimental for $K=4$. Individually tuning the noise levels for grids and particles can modestly lower error rates, but is either very sensitive or negligible. The model performs better when bundling more timesteps, enabling faster rollouts as a single forward pass predicts more steps. We found $m=8$ to be optimal with the other default hyperparameters, outperforming larger bundling. This is because more network capacity is needed to extract information for the next $16$ or $32$ timesteps from a single state. Instead, we opted for a shallower and narrower network to balance speed and memory footprint with performance gains. In terms of network architecture, we chose a U-Net. We experiment with an FNO [51] in Appendix B and find it to underperform, leading us to keep the U-Net architecture. We find the U-Net’s width and depth to have a minor impact on performance, confirming that a larger network is not needed. The grid size, however, is critical. A low resolution loses fine details, while a high resolution turns meaningful structures, such as liquid blobs or walls, into isolated voxels.
Given these results, we conduct the rest of our experiments using the blue parameters in Figure 3, except for WBC-SPH. As that dataset contains more particles and has a much longer rollout, we opted for bundling more steps ($m=16$), coupled with a more expressive depth of $4$. Crucially, as the water density is higher in that dataset, we increase the grid size to $128$ to capture finer interactions.
@@ 4.2 Comparison with previous work
We compare NeuralMPM against GNS [18] and DMCF [20]. We use the official implementations and training instructions released by the authors to assess training times, inference times, as well as accuracy. We compare against both GNS and DMCF on WaterRamps, Sandramps, and Goop. We also compare against GNS on MultiMaterial, and against DMCF on WBC-SPH. Additionally, we compare against the GNS and SEGNN [56] baselines provided by LagrangeBench (Appendix C).
@@
We report quantitative results comparing the long-term accuracy in Table 1 and show trajectories of NeuralMPM in Figure 2, as well as comparisons against baselines on WaterRamps in Figure 4. On the mono-material datasets WaterRamps, SandRamps, and Goop, NeuralMPM performs competitively with GNS and better than DMCF in terms of mean squared error (MSE). For MultiMaterial, NeuralMPM reduces the MSE by almost half, which we attribute to it being a hybrid method, known to better handle interactions, mixing, and collisions between different materials. In terms of Earth Mover’s Distance (EMD), NeuralMPM outperforms both baselines, suggesting that NeuralMPM is better at capturing the spatial distribution of the particles. However, on WBC-SPH, NeuralMPM falls behind DMCF. This dataset is challenging as SPH solvers outperform MPM solvers in this domain. It is also limited to only 30 training simulations, compared to 1000 in others, with trajectories reaching a steady state after 1000 timesteps (out of 3200), strongly reducing the quantity of data with information about the dynamics. The dataset also features variable gravity across simulations, which DMCF directly integrates as an external force in its update step. This approach could improve NeuralMPM’s performance on this dataset by offloading learning to apply gravity from the model.
Training.
Inference time.
Table 2 reports the average inference time for one model call and a full rollout on WaterRamps. NeuralMPM, despite predicting multiple frames per model call, achieves faster inference time per call than the baselines while maintaining comparable accuracy. Time bundling results in an even larger gap for full rollouts. Additionally, our method’s improved memory efficiency allows for over 100 parallel simulations, whereas GNS and DMCF face memory limitations with more than a few.
@@
Ground Truth
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ours
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
GNS
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
DMCF
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4: Example WaterRamps trajectory against baselines. We select a random test trajectory that was not seen during training and unroll predictions from NeuralMPM and the two baselines, starting from the same initial conditions. We display six snapshots, spaced evenly in time over the 600-step sequence.
$1$ Hour $5$ Hours $15$ Hours Best Ground Truth
Ours
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
GNS
Refer to caption Refer to caption Refer to caption Refer to caption
DMCF
Refer to caption Refer to caption Refer to caption Refer to caption
@@
NeuralMPM (Ours) GNS DMCF
One model call ($T=1$) 8.04 $\pm$ 0.43 13.91 $\pm$ 2.83 12.85 $\pm$ 8.41
Rollout ($T=600$) 606.63 $\pm$ 7.69 8344.94 $\pm$ 1700.8 9448.17 $\pm$ 804.07
Table 2: Inference time (in ms) of NeuralMPM and baselines on WaterRamps. Times were averaged over all validation trajectories. NeuralMPM predicts 8 frames in a single model call and still outperforms the two baselines per call, which further widens the gap for the total rollout time. For comparison, [18] reports a computation time of 71ms to simulate one timestep with Taichi-MPM.
@@ 4.3 Generalization
One notable advantage of NeuralMPM is its invariance to the number of particles, as the transition model only processes the voxelized representation. To demonstrate this, we train a model on WaterRamps, which contains about $2.3$k particles and $600$ timesteps, and evaluate it on WaterDrop-XL, which features about four times more particles, $1000$ timesteps, and no obstacles. An example snapshot is displayed in Figure 6. Interestingly, the larger number of particles only affects interpolation steps between the grid and particles, resulting in a negligible impact on total inference time, making the model nearly as fast despite 4 times more particles. We also validate generalization quantitatively by comparing the error rates on WaterDrop-XL of a model trained directly on it and the model trained solely on WaterRamps. With the same training budget, the latter achieves a lower MSE at $20.92\times 10^{-3}$ against $28.09\times 10^{-3}$. More trajectories are displayed in Figure 10.
If a domain-agnostic processor architecture is used, such as a fully convolutional U-Net or an FNO, then NeuralMPM can generalize to different domain shapes without retraining, as shown in Fig 6. We demonstrate this ability by considering a model solely trained on WaterRamps, a square domain of size $0.84\times 0.84$ mapped to $64\times 64$ grids. Without retraining, we perform inference with this model on larger unseen environments of size $1.68\times 0.84$, and change the grid size to $128\times 64$. The unseen environments were built by merging and modifying initial conditions of held-out test trajectories from WaterRamps. NeuralMPM manages to emulate particles in this larger and rectangular domain flawlessly despite being trained on a smaller domain with a smaller grid, showcasing the advantage of a U-Net for handling more general domains. No ground truth is displayed as [18] have not provided any information for simulating their data using Taichi-MPM. More trajectories are shown in Figure 11.
WaterDrop-XL Larger rectangular domains
Initial
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Initial
Predictions
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Predictions $t_{1}$
Ground Truth
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Predictions $t_{2}$
@@ 4.4 Inverse design problem
Test 1 Test 2 Test 3
Initial
@@
Unoptimized
Refer to caption Refer to caption Refer to caption
Optimized
Refer to caption Refer to caption Refer to caption
@@
Finally, we demonstrate the application of NeuralMPM for inverse problems on a toy inverse design task that consists in optimizing the direction of a ramp to make the particles reach a target location, similar to [4]. We place a blob of water at different starting locations, and we then place a ramp at some location, with a random initial angle $\alpha$. The goal is to spin that ramp by tuning $\alpha$ in order to make the water end up at a desired location. The main challenges of this task are the long-range time horizon of the goal and the presence of nonlinear physical dynamics. We proceed by selecting the point where we want the water to end up and compute the average distance between the point and particles at the last simulation frame (Tests 1 & 2) or at an intermediate frame (Test 3). We then minimize the distance via gradient descent, leveraging the differentiability of NeuralMPM to solve this inverse design problem. We show three example optimizations in Figure 7.
@@ 5 Conclusion
Like other approaches, NeuralMPM is limited by the computation used to process the structure of the point cloud. In our case, voxelization means we cannot deal with particles that lie outside of the domain and are limited to regular grids. Additionally, the size of the voxels is directly related to the fluid’s density, defined as the number of particles within a given volume. If the voxels are too large, the model will fail to capture finer details. Conversely, if they are too small, the model may struggle due to insufficient local structure. Similarly, performance can degrade in very sparse domains. Compressible fluids might also present challenges, though this requires further verification.
@@
Our work is only a first step towards hybrid Eulerian-Lagrangian neural emulators, leaving many avenues for future research. Extending NeuralMPM to 3D systems is a natural continuation of this work. Future studies could also explore alternative particle-to-grid and grid-to-particle functions, like the non-uniform Fourier transform [57], or more sophisticated interpolation methods from classical MPM literature [22]. Additionally, incorporating features commonly used in classical SPH and MPM simulations, such as viscosity, pressure, or temperature, presents another promising direction for future work. A less traditional direction is to make NeuralMPM probabilistic and encode richer distributional information about the particles in the grid nodes, instead of maintaining only a mean value. This could potentially improve NeuralMPM’s ability to resolve subgrid phenomena. Advances in Lagrangian Particle Tracking [58] will eventually make it possible to create datasets from real-world data, enabling the training of NeuralMPM directly from data without the need for the costly design process of a numerical simulator. Lastly, losses that take into account the distribution of particles and are invariant under particle permutation, such as the earth mover’s distance, are increasingly becoming an option to consider, as more efficient differentiable approximations emerge.

### v2

@@ 4 Experiments
We conduct a series of experiments to demonstrate the accuracy, speed, and generalization capabilities of NeuralMPM. Specifically, we examine its robustness to hyperparameter and architectural choices through an ablation study (4.1). We compare NeuralMPM to GNS and DMCF in terms of accuracy, training time, convergence, and inference speed (4.2). We also evaluate the generalization capabilities of NeuralMPM (4.3) and illustrate how its differentiability can be leveraged to solve an inverse design problem (4.4). Through these experiments, we demonstrate that NeuralMPM is a flexible, accurate, and fast method for emulating complex particle-based simulations. The baselines established by Winchenbach & Thuerey (2024) and hybrid simulators (Li et al., 2024, Ma et al., 2023) have promising results. However, we do not compare against them as they either use different benchmarks or are specifically tailored for certain physical domains, requiring material-specific knowledge. In contrast, NeuralMPM, like GNS and DMCF, requires only particle positions without being restricted to any particular domain.
@@
We consider 6 datasets with variable sequence lengths, numbers of particles, and materials. The first three datasets, WaterRamps, SandRamps, and Goop, contain a single material, water, sand, and goop, respectively, with different material properties. The first two datasets contain random ramp obstacles to challenge the model’s generalization capacity. The fourth dataset, MultiMaterial, mixes the three materials together in the same simulations. These four datasets are taken from Sanchez-Gonzalez et al. (2020) and were simulated using the Taichi-MPM simulator (Hu et al., 2018b). They each contain 1000 trajectories for training and 30 (Goop) or 100 (WaterRamps, SandRamps, MultiMaterial) for validation and testing. The fifth dataset, Dam Break 2D, was generated using SPH and contains 50 trajectories for learning, and 25 for validation and testing. The last dataset, VariableGravity, was also generated using Taichi-MPM. It consists of simulations with variable gravity of a water-like material, and contains 1000 trajectories for training and 100 for validation and testing.
@@
NeuralMPM is trained on trajectories with varying initial conditions and number of particles. The training batches are sampled randomly in time and across sequences. We use Adam (Kingma & Ba, 2014) with the following learning rate schedule: a linear warm-up over 100 steps from $10^{-5}$ to $10^{-3}$, $900$ steps at $10^{-3}$, then a cosine annealing (Loshchilov & Hutter, 2017) for $100,000$ iterations. We use a batch size of $128$, $K=4$ autoregressive steps per iteration, bundle $m=8$ timesteps per model call (resulting in $24$ predicted states), and a grid size of $64\times 64$. For most of our experiments, we use a U-Net (Ronneberger et al., 2015) with three downsampling blocks with a factor of $2$, $64$ hidden channels, a kernel size of $3$, and MLPs with three hidden layers of size 66664 for pixel-wise encoding and decoding into a latent space. For a fair comparison, we ran training and inference for NeuralMPM, DMCF, and GNS on the exact same hardware. GNS and DMCF were trained until convergence (a maximum of 120 and 240 hours, respectively), while NeuralMPM required 20 hours or less to converge. For WaterRamps, Sandramps, Goop, and MultiMaterial, we use the same parameters as those reported by authors. We hyperparameter search DMCF for Dam Break 2D and both GNS and DMCF for VariableGravity and report the best performance obtained for a budget of 60 GPU-days per dataset. Further details on training can be found in Appendix A.
@@
Figure 2: Example snapshots. We train and evaluate NeuralMPM on WaterRamps, SandRamps and Goop, each consisting of a single material, on MultiMaterial that mixes water, sand and goop, and on Dam Break 2D, a rectangular-shaped SPH dataset. NeuralMPM is able to learn various kinds of materials, their interactions, and their interactions with solid obstacles. Despite being inspired by MPM, it is not limited to data showing MPM-like behaviour.
@@
Figure 3: Ablation results. Mean squared error (MSE) of full rollouts on unseen test data for Goop. The default parameters are in blue. The dashed orange line shows the best rollout MSE of GNS as reported by Sanchez-Gonzalez et al. (2020) on the same data ($1.89\times 10^{-3}$), while the dotted line ($2.4\times 10^{-3}$) indicates the MSE we obtained for GNS after $240$ hours ($20$M training steps). NeuralMPM is robust to hyperparameter changes, with the biggest effects coming from the number of timesteps bundled together ($m$) and grid noise. For a rollout of length $T$, the model is called $T/m$ times, meaning lower values of $m$ require maintaining stability for longer. Autoregressive training coupled with time bundling is sufficient to stabilize the model, eliminating the need for noise injection. Although GNS reportedly outperforms NeuralMPM by a small margin, these results could not be reproduced in our experiments.
Figure 4: Example WaterRamps trajectory against baselines. Each method is unrolled starting from the initial conditions of a random test trajectory not seen during training. We display six snapshots evenly spaced in time.
@@ 4.1 Ablation study
Figure 3 summarizes the ablation results. A larger number $K$ of autoregressive steps yields more accurate rollouts without the need to add noise. Indeed, injecting noise does not improve accuracy and is even detrimental for $K=4$. Individually tuning the noise levels for grids and particles can modestly lower error rates, but is either very sensitive or negligible. The model performs better when bundling more timesteps, enabling faster rollouts as a single forward pass predicts more steps. We found $m=8$ to be optimal with the other default hyperparameters, outperforming larger bundling. This is because more network capacity is needed to extract information for the next $16$ or $32$ timesteps from a single state. Instead, we opted for a shallower and narrower network to balance speed and memory footprint with performance gains. In terms of network architecture, we chose a U-Net. We experiment with an FNO (Li et al., 2021) in Appendix B and find it to underperform, leading us to keep the U-Net architecture. We find the U-Net’s width and depth to have a minor impact on performance, confirming that a larger network is not needed. The grid size, however, is critical. A low resolution loses fine details, while a high resolution turns meaningful structures, such as liquid blobs or walls, into isolated voxels.
@@ 4.2 Comparison with previous work
We compare NeuralMPM against GNS and DMCF. We use the official implementations and training instructions to assess training times, inference times, as well as accuracy. We compare against both GNS and DMCF on WaterRamps, Sandramps, Goop, Dam Break 2D, and VariabelGravity. We also compare against GNS on MultiMaterial, but not against DMCF since it does not support multiple materials.
@@
We report quantitative results comparing the long-term accuracy in Table 1 and show trajectories of NeuralMPM in Figure 2, as well as comparisons against baselines on WaterRamps in Figure 4. On the mono-material datasets WaterRamps, SandRamps, and Goop, NeuralMPM performs competitively with GNS and better than DMCF in terms of mean squared error (MSE). For MultiMaterial, NeuralMPM reduces the MSE by almost half, which we attribute to it being a hybrid method, known to better handle interactions, mixing, and collisions between different materials. In Dam Break 2D, NeuralMPM outperforms both baselines, despite the data being simulated using SPH. Finally, NeuralMPM surpasses the performance of DMCF in VariableGravity, even though the latter accounts for gravity explicitly. In terms of Earth Mover’s Distance (EMD), NeuralMPM outperforms both baselines across all benchmarks, suggesting that NeuralMPM is better at capturing the spatial distribution of the particles.
@@
Figure 6: Time and memory performance. Average FPS (left) and GPU VRAM usage (right) for increasing numbers of particles for a traditional solver (Taichi-MPM (Hu et al., 2018a)), NeuralMPM, and the two baselines. The two baselines quickly require very large amounts of memory and become very slow. Although Taichi-MPM is more memory efficient for high numbers of particles, NeuralMPM remains much faster, emulating 30 million particles at 25FPS.
Training.
Inference time and memory.
In Figure 6, we display the time and memory performance of NeuralMPM, the two baselines GNS and DMCF, and the reference solver Taichi-MPM. In terms of speed, NeuralMPM strongly outperforms all three methods, partly thanks to time bundling, which considerably reduces the number of model calls required for a given number of frames to emulate. In terms of memory, although NeuralMPM remains inferior to Taichi-MPM, which is highly optimized, it can emulate tens of millions of particles on a single GPU, while GNS and DMCF struggle to reach half a million.
@@ 4.3 Generalization
One notable advantage of NeuralMPM is its invariance to the number of particles, as the transition model only processes the voxelized representation. To demonstrate this, we train a model on WaterRamps, which contains about $2.3$k particles and $600$ timesteps, and evaluate it on WaterDrop-XL, which features about four times more particles, $1000$ timesteps, and no obstacles. An example snapshot is displayed in Figure 7. The larger number of particles only affects interpolation steps between the grid and particles, resulting in a negligible impact on total inference time, making the model nearly as fast despite 4 times more particles. We also validate generalization quantitatively by comparing the error rates on WaterDrop-XL of a model trained directly on it and the model trained solely on WaterRamps. With the same training budget, the latter achieves a lower MSE at $20.92\times 10^{-3}$ against $28.09\times 10^{-3}$. More trajectories are displayed in Figure 18.
If a domain-agnostic processor architecture is used, such as a fully convolutional U-Net or an FNO, then NeuralMPM can generalize to different domain shapes without retraining, as shown in Fig 7. We demonstrate this ability by considering a model solely trained on WaterRamps, a square domain of size $0.84\times 0.84$ mapped to $64\times 64$ grids. Without retraining, we perform inference with this model on larger unseen environments of size $1.68\times 0.84$, and change the grid size to $128\times 64$. The unseen environments were built by merging and modifying initial conditions of held-out test trajectories from WaterRamps. NeuralMPM emulates particles in this larger and rectangular domain despite being trained on a smaller square domain with a smaller grid, showing that a U-Net can generalize to other domains. No ground truth is displayed as Sanchez-Gonzalez et al. (2020) provide no information about the data generation. More trajectories are shown in Figure 19.
@@ 4.4 Inverse design problem
Finally, we demonstrate the application of NeuralMPM for inverse problems on a toy inverse design task that consists in optimizing the direction of a ramp to make the particles reach a target location, similar to (Allen et al., 2022). We place a blob of water at different starting locations, and we then place a ramp at some location, with a random initial angle $\alpha$. The goal is to spin the ramp by tuning $\alpha$ in order to make the water end up at a desired location. The main challenges of this task are the long-range time horizon of the goal and the presence of nonlinear physical dynamics. We proceed by selecting the point where we want the water to end up and compute the average distance between the point and particles at the last simulation frame. We then minimize the distance via gradient descent, leveraging the differentiability of NeuralMPM to solve this inverse design problem. We show am example optimization in Figure 8, and additional examples in Appendix B.
Initial Unoptimized Optimized
@@ 5 Conclusion
Like other approaches, NeuralMPM is limited by the computation used to process the structure of the point cloud. In our case, voxelization means we cannot deal with particles that lie outside of the domain and are limited to regular grids. Additionally, the size of the voxels is directly related to the number of particles within a given volume. If the voxels are too large, the model will fail to capture finer details. Conversely, if they are too small, the model may struggle due to insufficient local structure. Similarly, performance can degrade in very sparse domains. Compressible fluids might also present challenges, though this requires further verification.
@@
Our work is only a first step towards hybrid Eulerian-Lagrangian neural emulators, leaving many avenues for future research. Extending NeuralMPM to 3D systems is a natural continuation of this work. Future studies could also explore alternative particle-to-grid and grid-to-particle functions, like the non-uniform Fourier transform (Fessler & Sutton, 2003), or more sophisticated interpolation methods from classical MPM literature (Nguyen et al., 2023). A less traditional direction is to make NeuralMPM probabilistic and encode richer distributional information about the particles in the grid nodes, instead of maintaining only a mean value. This could potentially improve NeuralMPM’s ability to resolve subgrid phenomena. Finally, advances in Lagrangian Particle Tracking (Schröder & Schanz, 2023) will eventually make it possible to create datasets from real-world data, enabling the training of NeuralMPM directly from data without the need for the costly design process of a numerical simulator.